# Neural Spectro-polarimetric Fields

### v1

@@ Abstract.
Modeling the spatial radiance distribution of light rays in a scene has been extensively explored for applications, including view synthesis. Spectrum and polarization — the wave properties of light — are often neglected due to their integration into three RGB spectral bands and their non-perceptibility to human vision. Despite this, these properties encompass substantial material and geometric information about a scene. In this work, we propose to model spectro-polarimetric fields, the spatial Stokes-vector distribution of any light ray at an arbitrary wavelength. We present Neural Spectro-polarimetric Fields (NeSpoF) — a neural representation that models the physically-valid Stokes vector at given continuous variables of position, direction, and wavelength. NeSpoF manages inherently noisy raw measurements, showcases memory efficiency, and preserves physically vital signals — factors that are crucial for representing the high-dimensional signal of a spectro-polarimetric field. To validate NeSpoF, we introduce the first multi-view hyperspectral-polarimetric image dataset, comprised of both synthetic and real-world scenes. These were captured using our compact hyperspectral-polarimetric imaging system, which has been calibrated for robustness against system imperfections. We demonstrate the capabilities of NeSpoF on diverse scenes.
Figure 1. We present neural spectro-polarimetric fields that allow for the rendering of a scene from novel views, not limited to the synthesis of (a) RGB radiance images. We can render (b) hyperspectral Stokes vectors for each individual light ray, thus providing a comprehensive description of its spectro-polarization property. Also, (c) it can inherently reduce noise in low-signal-to-noise ratio spectro-polarimetric measurements, as demonstrated in the degree of polarization (DoP). We recommend viewing this with a PDF reader for clarity. (d) We further show the analysis of spectro-polarimetric information for the CD that is otherwise invisible to the human eye, as illustrated by the radiance and angle-of-linear-polarization.
@@ 1. Introduction
With the objective of harnessing multiple dimensions of light concurrently, recent methods have utilized both polarization and spatial dimensions. These have found applications such as 3D reconstruction with a polarized light stage [Ghosh et al., 2011] and multi-view capture with a hand-held polarization camera [Hwang et al., 2022; Dave et al., 2022]. Image segmentation leveraging polarization, spectrum, and spatial dimensions was demonstrated by Chen et al. [2022]. However, the exploitation of high-dimensional data still remains challenging.
Specifically, the representation of such high-dimensional visual information should facilitate physically-valid modeling of light properties without using excessive computational resources. Tensor representation [Baek and Heide, 2021; Chen et al., 2022] has commonly been used to handle partial dimensions of light properties, assigning each light property to a distinct dimension of the tensor. However, the joint modeling of spectrum, polarization, position, and direction can lead to prohibitive memory requirements, potentially escalating to 58 GB [Karnewar et al., 2022]. Also, the interpolation of polarization properties for unseen position, direction, and wavelength can result in physically invalid polarization representation. These representations are also prone to measurement noise, which can easily arise due to the cumulative effect of spectral and polarization filters [Chen et al., 2022], and a low pixel count [Fan et al., 2022].
In this work, we propose to model spectro-polarimetric fields that describe the Stokes vector of a light ray, as a polarimetric representation [Collett, 2005; Wilkie and Weidlich, 2010], at a position, direction, and wavelength. We present Neural Spectro-polarimetric Fields, NeSpoF, the first neural representation specifically designed to represent physically-valid spectro-polarimetric fields. In addition to the modeling capability of physically-valid Stokes vectors, NeSpoF benefits from the regularization capability of implicit neural networks, enabling it to handle low-SNR measurements. Also, NeSpoF has a compact memory footprint of 13.6 MB compared to the tensor-based model of 58 GB. Refer to the Supplemental Document for the memory analysis. To evaluate NeSpoF on real-world scenes, we implement a hyperspectral-polarimetric imaging system and its multi-view capture procedure, for which we propose a calibration method to account for the physical imperfections of the acquisition system. Using the imaging system, we capture and release the first multi-view hyperspectral-polarimetric image datasets. We demonstrate the view-spectrum-polarization synthesis using NeSpoF.
@@ 2. Related Work
2.1. Representation for Multi-dimensional Light Properties
@@
Computational representation of multi-dimensional light properties is a critical consideration across visual computing applications. A prevalent method is the tensor-based approach, where each tensor dimension corresponds to a distinct light property. This approach spans the spatial tensor with two-plane methods and lumigraphs [Levoy and Hanrahan, 1996; Gortler et al., 1996; Sitzmann et al., 2019], hyperspectral images [Kim et al., 2012], and multi-dimensional tensors [Baek and Heide, 2021; Chen et al., 2022]. However, the tensor representation for position, direction, spectrum, and polarization can impose an excessive memory burden due to the increased dimensionality. Moreover, it is vulnerable to the low SNR of hyperspectral-polarimetric measurements [Chen et al., 2022].
@@
Recently, coordinate-based neural networks modeling the spatial distribution of RGB light intensity have exhibited promising performance in novel-view synthesis [Mildenhall et al., 2020; Suhail et al., 2022; Mildenhall et al., 2022]. PANDORA [Dave et al., 2022] utilizes polarization for robust neural 3D reconstruction of specular surfaces via a polarimetric BRDF model [Baek et al., 2018]. However, it does not directly model the spectro-polarimetric field of Stokes vectors under variable wavelengths, and it relies on the limited representational capacity of the polarimetric BRDF model. pCON [Peters et al., 2023] represents a polarization image based on singular-value-decomposition via a sinusoid-activation neural network, which considers frequency differences between intensity and polarimetric properties. However, it does not model the directional dimension and is restricted to trichromatic channels. In this work, for the first time, we propose to model a physically-valid spectro-polarimetric field using NeSpoF.
2.2. Applications using Multi-dimensional Light Properties
@@
The individual dimensions of light properties have been used in various applications. The position and direction have been utilized for view synthesis and geometry reconstruction [Penner and Zhang, 2017; Mildenhall et al., 2020; Guo et al., 2022; Mildenhall et al., 2019; Jin et al., 2005; Liu et al., 2022]. The spectral dimension has been used for detection and segmentation [Imamoglu et al., 2018; Wang et al., 2021; Trajanovski et al., 2020; Aloupogianni et al., 2022] and for estimating illumination spectrum [Li et al., 2021; An et al., 2015]. Polarization has been used for shape reconstruction [Kadambi et al., 2015; Lei et al., 2022; Zou et al., 2020; Ba et al., 2020; Baek et al., 2018; Ding et al., 2021; Fukao et al., 2021], acquiring appearance [Deschaintre et al., 2021; Kondo et al., 2020], removing reflections [Nayar et al., 1997; Wen et al., 2021; Yang et al., 2016], segmenting transparent objects [Kalra et al., 2020; Mei et al., 2022], seeing through scattering [Zhou et al., 2021; Baek and Heide, 2021; Liu et al., 2015], and tone mapping [del Molino and Muñoz, 2019]. Using multiple light properties could further enhance these capabilities such as shape and appearance reconstruction with the spatial and polarimetric dimensions [Baek and Heide, 2021; Dave et al., 2022; Riviere et al., 2017]. Spectral-polarimetric analysis has been used for object recognition [Denes et al., 1998], skin analysis [Zhao et al., 2009], and dehazing [Xia and Liu, 2016]. In this work, using NeSpoF, we demonstrate view-spectrum-polarization synthesis, which reveals various hidden scene properties.
2.3. Acquisition of Multi-dimensional Light Properties
@@
Several methods have been proposed for simultaneously acquiring multiple dimensions of light. The positional and directional dimensions can be acquired using hand-held multi-view capture of images [Davis et al., 2012; Mildenhall et al., 2020, 2019] or multi-view camera arrays [Yang et al., 2002; Wilburn et al., 2004, 2005]. Capturing the positional, directional, and polarimetric dimensions has been practiced with light stages equipped with polarizers [Ghosh et al., 2011], hand-held polarization cameras and flashlights [Hwang et al., 2022], and polarization cameras mounted on goniometers [Dave et al., 2022]. Light-field hyperspectral imaging systems [Cui et al., 2021; Manakov et al., 2013] and hyperspectral camera with a rotation stage [Kim et al., 2012] allow for capturing the directional and spectral dimensions. Hyperspectral-polarimetric imaging allows for capturing the spectral and polarimetric dimensions [Denes et al., 1998; Fan et al., 2020, 2022; Altaqui et al., 2021; Bai et al., 2021; Chen et al., 2022]. To test NeSpoF, we implement a scanning-based hyperspectral-polarimetric imaging system and propose calibrating the imperfections of the imaging system. Using the hyperspectral-polarimetric camera, we release the first real-world dataset of multi-view hyperspectral-polarimetric images that can be used for training NeSpoF and may spur further interest in multi-dimensional visual analysis.
@@ 3.1. Neural Radiance Fields
The plenoptic function $f^{p}(\cdot)$ is a multi-dimensional function that describes the spatial and temporal distribution of light in a 3D space by modeling the light intensity $L$ as a function of a 3D position $(x,y,z)$, viewing direction $(\theta,\phi)$, wavelength $\lambda$, and time $\tau$ [Adelson and Wang, 1992]: $L=f^{p}(x,y,z,\theta,\phi,\lambda,\tau)$. Assuming static scenes and constant illumination allows for detaching the temporal dependency, resulting in the static plenoptic function: $L=f^{p}(x,y,z,\theta,\phi,\lambda)$.
Neural Raidance Field (NeRF) [Mildenhall et al., 2020] represents a simplified plenoptic function and density of a scene as a continuous volumetric field modeled with an MLP $F_{\Theta}^{\text{NeRF}}$ as
@@
where $\mathbf{L}=[L_{R},L_{G},L_{B}]^{\intercal}$ is the RGB radiance at the position $(x,y,z)$ sampled from the direction $(\theta,\phi)$. $\sigma$ is the volume density at the position $(x,y,z)$. $\Theta$ is the network weights. Figure 3(a) depicts the RGB radiance of a ray. Rendering the RGB values of a pixel at the position $\mathbf{p}$ amounts to casting rays and volumetrically integrating the learned radiance $\mathbf{L}$ and density $\sigma$.
@@ 3.2. Polarization
Polarization is a wave property of light that describes the oscillation pattern of the electric field of light [Wilkie and Weidlich, 2010; Collett, 2005]. The Stokes-Mueller formalism is a mathematical framework that describes the polarization state of light and how it changes as it travels through a medium. The Stokes vector, denoted by $\mathbf{s}=[s_{0},s_{1},s_{2},s_{3}]^{\intercal}$, describes the complete polarization state of light. The four elements of the vector represent different properties of the light. $s_{0}$ represents the total radiance of the light. $s_{1}$ and $s_{2}$ represent the differences in the radiance of linearly-polarized light at ${}^{\circ}$${}^{\circ}$ and ${}^{\circ}$${}^{\circ}$ respectively. $s_{3}$ represents the difference in the radiance of right- and left-circularly polarized light. For instance, horizontal linearly-polarized light can be modeled as a Stokes vector $\mathbf{s}=[1,1,0,0]^{\intercal}$. The Stokes vector $\mathbf{s}$ is physically valid if the elements satisfy the condition $s_{0}^{2}\geq s_{1}^{2}+s_{2}^{2}+s_{3}^{2}$. The Mueller matrix $\mathbf{M}\in\mathbb{R}^{4\times 4}$ describes the change of the polarization state of a light ray. To represent the change of a Stokes vector, the Mueller matrix is multiplied to a Stokes vector $\mathbf{s}_{\text{in}}$: $\mathbf{s}_{\text{out}}=\mathbf{M}\mathbf{s}_{\text{in}}$, where $\mathbf{s}_{\text{out}}$ is the output Stokes vector.
@@
A Stokes vector is associated with a coordinate system $\{\mathbf{x},\mathbf{y},\mathbf{z}\}$, where $\mathbf{z}$ represents the direction of light propagation. The $\mathbf{x}$ and $\mathbf{y}$ can be chosen arbitrarily as long as they form an orthonormal coordinate system with $\mathbf{z}$, as shown in Figure 2(a). This means that the polarization state of a light ray can be described differently with Stokes vectors with different elements lying at different coordinates. For instance, the Stokes vector $[1,1,0,0]^{\intercal}$ defined in the coordinate system of $\{\mathbf{x},\mathbf{y},\mathbf{z}\}$ describes the same polarization state with the Stokes vector $[1,-1,0,0]^{\intercal}$ in the coordinate system of $\{\mathbf{R}^{90^{\circ}}_{\mathbf{z}}\mathbf{x},\mathbf{R}^{90^{\circ}}_{\mathbf{z}}\mathbf{y},\mathbf{R}^{90^{\circ}}_{\mathbf{z}}\mathbf{z}\}$, where $\mathbf{R}^{90^{\circ}}_{\mathbf{z}}$ is a rotation matrix by 90 degrees about the $\mathbf{z}$ axis.
Figure 2. (a) The Stokes vector is defined on a coordinate system $\{\mathbf{x},\mathbf{y},\mathbf{z}\}$, where $\mathbf{x},\mathbf{y}$ can be set arbitrary to form an orthonormal coordinate with the propagation direction $\mathbf{z}$. (b) The polarization state of light can be described as a projection of the oscillating electric field onto the tangent plane, which is called polarization ellipse, where $\psi$ and $\chi$ are the azimuth and ellipticity angles. (c) The Poincaré sphere visualizes the Stokes vector in 3D with the normalized Stokes-vector elements.
Figure 3. (a) NeRF models a simplified plenoptic function to represent spatially-distributed RGB radiance. (b) Spectro-polarimetric field describes the polarization state and intensity of light as a Stokes vector per each wavelength $\lambda$, position $x,y,z$, and direction $\theta,\phi$. (c) NeSpoF, consisting of a positional MLP $F_{\Theta}^{p}$ and a spectro-directional MLP $F_{\Theta}^{d}$, is a neural representation that describes the polarization state and radiance of spatially-distributed light at each wavelength. The inputs to the network are the wavelength $\lambda$, position $(x,y,z)$ and direction $\theta,\phi$, which are both defined on the world coordinate system $\{\mathbf{x}_{w},\mathbf{y}_{w},\mathbf{z}_{w}\}$. We apply positional encoding $\gamma_{k}$ of the frequency $k$ for each dimension. The outputs of the network are the volume density $\sigma$ at the position $(x,y,z)$ and polarization parameters ($s_{0}$, $\rho$, $\chi$, and $\psi$). The polarization parameters are converted to the physically-valid Stokes vector $\mathbf{s}$, defined on the Stokes-vector coordinate system $\{\mathbf{x}_{s}^{w},\mathbf{y}_{s}^{w},\mathbf{z}_{s}^{w}\}$.
@@
Visualization of Stokes Vector

### v2

@@ Abstract.
Modeling the spatial radiance distribution of light rays in a scene has been extensively explored for applications, including view synthesis. Spectrum and polarization, the wave properties of light, are often neglected due to their integration into three RGB spectral bands and their non-perceptibility to human vision. However, these properties are known to encompass substantial material and geometric information about a scene. Here, we propose to model spectro-polarimetric fields, the spatial Stokes-vector distribution of any light ray at an arbitrary wavelength. We present Neural Spectro-polarimetric Fields (NeSpoF), a neural representation that models the physically-valid Stokes vector at given continuous variables of position, direction, and wavelength. NeSpoF manages inherently noisy raw measurements, showcases memory efficiency, and preserves physically vital signals — factors that are crucial for representing the high-dimensional signal of a spectro-polarimetric field. To validate NeSpoF, we introduce the first multi-view hyperspectral-polarimetric image dataset, comprised of both synthetic and real-world scenes. These were captured using our compact hyperspectral-polarimetric imaging system, which has been calibrated for robustness against system imperfections. We demonstrate the capabilities of NeSpoF on diverse scenes.
Neural fields, hyperspectral imaging, polarimetric imaging
View synthesis, hyperspectral imaging, polarimetric imaging
submissionid: 412journalyear: 2023copyright: rightsretainedconference: SIGGRAPH Asia 2023 Conference Papers; December 12–15, 2023; Sydney, NSW, Australiabooktitle: SIGGRAPH Asia 2023 Conference Papers (SA Conference Papers ’23), December 12–15, 2023, Sydney, NSW, Australiaprice: 15.00doi: 10.1145/3610548.3618172isbn: 979-8-4007-0315-7/23/12ccs: Computing methodologies Computer graphicsccs: Computing methodologies Computer vision representations
Figure 1. We present neural spectro-polarimetric fields that allow for the rendering of a scene from novel views, not limited to the synthesis of (a) RGB radiance images. We can render (b) hyperspectral Stokes vectors for each individual light ray, thus providing a comprehensive description of its spectro-polarization property. Also, (c) it can inherently reduce noise in low-signal-to-noise ratio spectro-polarimetric measurements, as shown in the degree of polarization (DoP). (d) We further show the analysis of spectro-polarimetric information for the CD that is otherwise invisible to the human eye.
@@ 1. Introduction
With the objective of harnessing multiple dimensions of light concurrently, recent methods have utilized both polarization and spatial dimensions. These have found applications such as 3D reconstruction with a polarized light stage [Ghosh et al., 2011] and multi-view capture with a hand-held polarization camera [Hwang et al., 2022; Dave et al., 2022]. Image segmentation leveraging polarization, spectrum, and spatial dimensions was also demonstrated by Chen et al. [2022].
While efforts have been directed towards applications, there remains a need for efficient and effective representations of high-dimensional light properties due to the curse of dimensionality affecting data size. For instance, tensor representations [Baek and Heide, 2021; Chen et al., 2022] have been used to handle partial dimensions of light properties, assigning each light property to a distinct dimension of the tensor. However, such direct modeling of spectrum, polarization, position, and direction can lead to prohibitive memory requirements. Moreover, these representations are prone to measurement noise, which can easily arise due to the cumulative effect of optical filters employed in spectral and polarimetric acquisition.
In this work, we propose to model spectro-polarimetric fields that describe the Stokes vector of a light ray, as a polarimetric representation [Collett, 2005; Wilkie and Weidlich, 2010], at a position, direction, and wavelength. As a representation for the spectro-polarimetric field, we present neural spectro-polarimetric fields (NeSpoF), the first neural representation specifically designed to represent physically-valid spectro-polarimetric fields. In addition to the physical plausibility of Stokes vectors, NeSpoF handles low-SNR measurements and exhibits a compact memory footprint of 13.6 MB, compared to a tensor-based model that consumes 58 GB (refer to the Supplemental Document). To evaluate NeSpoF on real-world scenes, we built a hyperspectral-polarimetric imaging system and its multi-view capture procedure, for which we propose a calibration method to account for the physical imperfections of the acquisition system. Using the imaging system, we capture and release the first multi-view hyperspectral-polarimetric image datasets. On the dataset, we demonstrate the synthesis capability of NeSpoF, taking into account the view, spectrum, and polarization.
Specifically, our main contributions are described as follows:
  • We take a first step on capturing, modeling, and rendering spectro-polarimetric fields and demonstrate learning view-spectro-polarimetric variations of real-world objects.
  • We develop a physically-valid neural representation, NeSpoF.
  • We propose the spatially-varying spectro-polarimetric calibration method to compensate the physical imperfections of the experimental prototype.
  • We release the first spectro-polarimetric multi-view dataset.
@@ 2. Related Work
2.1. Representations of Light Properties
@@
Computational representation of multi-dimensional light properties is a critical consideration across visual computing applications. A prevalent method is the tensor-based approach, where each tensor dimension corresponds to a distinct light property. This approach spans the spatial tensor with two-plane methods and lumigraphs [Levoy and Hanrahan, 1996; Gortler et al., 1996; Sitzmann et al., 2019], hyperspectral images [Kim et al., 2012], and multi-dimensional tensors [Baek and Heide, 2021; Chen et al., 2022]. However, the tensor representation for position, direction, spectrum, and polarization can impose an excessive memory burden due to the increased dimensionality. Moreover, it is vulnerable to the low SNR of typical hyperspectral-polarimetric measurements.
@@
Recently, coordinate-based neural representations [Mildenhall et al., 2020; Suhail et al., 2022; Mildenhall et al., 2022] and explicit primitives [Kerbl et al., 2023; Zhang et al., 2022] modeling the spatial distribution of RGB light intensity have emerged. Two related neural representations exist in the context of polarization and spatial dimension. PANDORA [Dave et al., 2022] utilizes polarization for robust neural 3D reconstruction of specular surfaces via a polarimetric BRDF model [Baek et al., 2018]. However, the polarimetric BRDF model limits the representational capacity of PANDORA, and it does not directly model Stokes vectors under variable wavelengths. pCON [Peters et al., 2023] represents a polarization image based on singular-value-decomposition via a sinusoid-activation neural network. It does not model the directional dimension and is restricted to trichromatic channels. In this work, we propose NeSpoF as a first neural representation for spectro-polarimetric fields.
2.2. Applications using Light Properties
@@
The individual properties of light have been used in various applications. View synthesis and geometry reconstruction often exploit the position and direction of light [Penner and Zhang, 2017; Mildenhall et al., 2020; Guo et al., 2022; Mildenhall et al., 2019; Jin et al., 2005; Liu et al., 2022]. The spectral dimension enables accurate detection and segmentation [Imamoglu et al., 2018; Wang et al., 2021; Trajanovski et al., 2020; Aloupogianni et al., 2022] and estimating illumination spectrum [Li et al., 2021; Khan et al., 2017; An et al., 2015]. Polarization provides information for shape reconstruction [Kadambi et al., 2015; Lei et al., 2022; Zou et al., 2020; Ba et al., 2020; Baek et al., 2018; Ding et al., 2021; Fukao et al., 2021], appearance acquisition [Deschaintre et al., 2021; Kondo et al., 2020], reflection removal [Nayar et al., 1997; Wen et al., 2021; Yang et al., 2016], transparent-object segmentation [Kalra et al., 2020; Mei et al., 2022], seeing through scattering [Zhou et al., 2021; Baek and Heide, 2021; Liu et al., 2015], and tone mapping [del Molino and Muñoz, 2019]. Simultaneously exploiting multiple light properties further enhances these capabilities, such as shape and appearance reconstruction with the spatial and polarimetric dimensions [Baek and Heide, 2021; Dave et al., 2022; Riviere et al., 2017]. Spectro-polarimetric analysis has been also used for object recognition [Denes et al., 1998], skin analysis [Zhao et al., 2009], and dehazing [Xia and Liu, 2016]. In this work, we demonstrate view-spectrum-polarization synthesis with NeSpoF, revealing complex spectro-polarimetric variations of real-world objects at novel views.
2.3. Acquisition of Light Properties
@@
Several methods have been proposed for simultaneously acquiring multiple dimensions of light. The positional and directional dimensions can be acquired using hand-held multi-view capture of images [Davis et al., 2012; Mildenhall et al., 2020, 2019] or multi-view camera arrays [Yang et al., 2002; Wilburn et al., 2004, 2005]. Capturing the positional, directional, and polarimetric dimensions has been practiced with light stages equipped with polarizers [Ghosh et al., 2011], hand-held polarization cameras and flashlights [Hwang et al., 2022], and polarization cameras mounted on goniometers [Dave et al., 2022]. Light-field hyperspectral imaging systems [Cui et al., 2021; Manakov et al., 2013], hyperspectral camera with a rotation stage [Kim et al., 2012], and metalens-based snapshot spectral light-field imaging [Hua et al., 2022] allow for capturing the directional and spectral dimensions. Hyperspectral-polarimetric imaging allows for capturing the spectral and polarimetric dimensions [Denes et al., 1998; Fan et al., 2020, 2022; Altaqui et al., 2021; Bai et al., 2021; Chen et al., 2022; Lv et al., 2020]. To test NeSpoF, we implement a scanning-based hyperspectral-polarimetric imaging system and propose calibrating the imperfections of the imaging system. Using the hyperspectral-polarimetric imaging system, we release the first real-world dataset of multi-view hyperspectral-polarimetric images that can be used for training NeSpoF and may spur further interest in multi-dimensional visual analysis.
@@ 3.1. Neural Radiance Fields
The plenoptic function $f^{p}(\cdot)$ is a multi-dimensional function that describes the spatial and temporal distribution of light in a 3D space by modeling the light intensity $L$ as a function of a 3D position $(x,y,z)$, viewing direction $(\theta,\phi)$, wavelength $\lambda$, and time $\tau$ [Adelson and Wang, 1992]: $L=f^{p}(x,y,z,\theta,\phi,\lambda,\tau)$. Assuming static scenes and constant illumination results in the static plenoptic function: $L=f^{p}(x,y,z,\theta,\phi,\lambda)$.
Neural radiance field (NeRF) [Mildenhall et al., 2020] represents the simplified plenoptic function and density of a scene as a continuous volumetric field parameterized with an MLP $F_{\Theta}^{\text{NeRF}}$ as
@@
where $\mathbf{L}=[L_{R},L_{G},L_{B}]^{\intercal}$ is the RGB radiance at the position $(x,y,z)$ sampled from the direction $(\theta,\phi)$. $\sigma$ is the volume density at the position $(x,y,z)$. $\Theta$ is the network weights. Rendering the RGB values of a pixel at the position $\mathbf{p}$ amounts to volumetrically integrating the learned radiance $\mathbf{L}$ and density $\sigma$.
@@ 3.2. Polarization
Polarization is a wave property of light that describes the oscillation pattern of the electric field of light [Wilkie and Weidlich, 2010; Collett, 2005]. The Stokes-Mueller formalism is a mathematical framework that describes the polarization state of light and how it changes. The Stokes vector, denoted by $\mathbf{s}=[s_{0},s_{1},s_{2},s_{3}]^{\intercal}$, describes the complete polarization state of light. The four elements of the vector represent different properties of the light. $s_{0}$ represents the total radiance of the light. $s_{1}$ and $s_{2}$ represent the differences in the radiance of linearly-polarized light at ${}^{\circ}$${}^{\circ}$ and ${}^{\circ}$${}^{\circ}$ respectively. $s_{3}$ represents the difference in the radiance of right- and left-circularly polarized light. The Mueller matrix $\mathbf{M}\in\mathbb{R}^{4\times 4}$ describes the change of the polarization state of a light ray by being multiplied to a Stokes vector $\mathbf{s}_{\text{in}}$: $\mathbf{s}_{\text{out}}=\mathbf{M}\mathbf{s}_{\text{in}}$, where $\mathbf{s}_{\text{out}}$ is the output Stokes vector.
@@
A Stokes vector is associated with a coordinate system $\{\mathbf{x},\mathbf{y},\mathbf{z}\}$, where $\mathbf{z}$ represents the direction of light propagation. The $\mathbf{x}$ and $\mathbf{y}$ can be chosen arbitrarily as long as they form an orthonormal coordinate system with $\mathbf{z}$, as shown in Figure 2(a). This means that the polarization state of a light ray can be described differently with Stokes vectors with different elements lying at different coordinates.
Figure 3. (a) NeRF models a simplified plenoptic function to represent spatially-distributed RGB radiance. (b) The spectro-polarimetric field describes the polarization state and intensity of light as a Stokes vector per each wavelength $\lambda$, position $x,y,z$, and direction $\theta,\phi$. (c) NeSpoF describes the spectro-polarimetric field, considering the physical validity of Stokes vectors and wavelength as an input variable. See Section 4 for details of NeSpoF.
@@
Visualization of Stokes Vectors